# Domain Generalization: A Tale of Two ERMs

### v1

@@ Abstract
Domain generalization (DG) is the problem of generalizing from several distributions (or domains), for which labeled training data are available, to a new test domain for which no labeled data is available. A common finding in the DG literature is that it is difficult to outperform empirical risk minimization (ERM) on the pooled training data.
In this work, we argue that this finding has primarily been reported for datasets satisfying a covariate shift assumption. When the dataset satisfies a posterior drift assumption instead, we show that “domain-informed ERM,” wherein feature vectors are augmented with domain-specific information, outperforms pooling ERM. These claims are supported by a theoretical framework and experiments on language and vision tasks.
fnum@section1 Introduction:
the ERM dilemma
@@
Domain generalization (DG) is a learning problem where the learner has access to labeled data from several source domains, and the goal is to generalize to a new target domain for which no labeled data is available. Let $X$ denote the input features, $Y$ the label, and $D$ the domain index.
@@
A persistent puzzle in DG is the surprising effectiveness of empirical risk minimization (ERM), a baseline that simply pools labeled data from all source domains together and trains a domain-agnostic classifier. Despite extensive efforts to design sophisticated DG algorithms, multiple studies have consistently shown that ERM remains highly competitive:
@@
Similar observations about the strong performance of ERM have been made on other DG benchmarks (Koh et al., 2021; Sagawa et al., 2022), as well as in related settings such as federated domain generalization (Bai et al., 2024).
A related observation is that most existing DG approaches learn a classifier that predicts $Y$ solely from $X$, thus ignoring domain information during inference. This is reflected in recent surveys:
  • Wang et al. (2022) (survey): “The goal of domain generalization is to learn a robust and generalizable predictive function $h:\mathcal{X}\to\mathcal{Y}$ from the $M$ training domains to achieve a minimum prediction error on an unseen test domain $S_{\text{test}}$.”
This is despite the fact that early works on DG learn predictions based not only on feature vectors, but also on domain-specific information (Blanchard et al., 2011; Muandet et al., 2013). While more recent methods also incorporate domain information into the training and inference procedure (Dubey et al., 2021; Bui et al., 2021; Yao et al., 2024), the fundamental role of domain information—its precise decision-theoretic value and the conditions under which it strictly improves performance—remains poorly understood.
In this work, we argue that the prevailing conclusions about ERM being “hard to beat” stem primarily from the fact that most benchmark DG datasets are from vision tasks. These datasets are characterized by a covariate shift assumption, which means that there exists a single classifier that performs well on all domains, and only the marginal distribution of $X$ changes from domain to domain. In such applications, strong performance is indeed possible without the use of domain-specific information, and ERM on pooled training data can perform quite well.
We consider DG problems characterized by posterior drift, where the conditional distribution of $Y|X$ (i.e., the posterior) changes with domain. We argue that for such DG problems, pooling ERM is inadequate, and stronger performance is achievable by “domain-informed” ERM, where domain-specific information is used both during training and at inference.
More broadly, a key goal of this work is to reassert (and extend) the importance of the original formulation of DG, which leverages domain-specific information when making predictions. We argue that it is this element of DG that makes it fundamentally distinct from conventional classification.
The contributions of this work are:
  • A theoretical framework extending the original formulation of DG by Blanchard et al. (2011).
  • Risk bounds that characterize when domain-specific information is beneficial (posterior drift) and when it is not (covariate shift).
  • A quantification of the difference between DG and domain adaptation, addressing an open question in Blanchard et al. (2021, Lemma 9).
  • Empirical validation of these findings on both language and vision tasks.
fnum@section2 Literature review
In these vision-based setups, the underlying distributional shift can be described as covariate shift (Ben-David et al., 2006; Mansour et al., 2009), where the marginal distribution $P_{X}$ varies significantly across domains, while the posterior $P_{Y|X}$ is constant. Importantly, domain information is frequently viewed as irrelevant or even spurious (Sagawa et al., 2020) for predicting labels. Consequently, much of the literature has focused on learning domain-invariant representations (Sun and Saenko, 2016; Ganin et al., 2016; Arjovsky et al., 2019). Additional references are in Appendix˜A.
In contrast, our work is motivated by a different class of problems characterized by posterior drift (Scott, 2019; Cai and Wei, 2021; Maity et al., 2024; Zhu et al., 2024; Wang and Ning, 2025), where the conditional distribution of $Y|X$ varies across domains. This type of shift commonly arises in natural language processing (NLP). For a given sentence $X$, different annotators—or populations—may interpret its semantic content differently, leading to divergent labels $Y$ (e.g., offensive vs. non-offensive, positive vs. negative). Such inherent ambiguity in language often results in systematic disagreement in annotations. These effects have been documented across a wide range of NLP tasks (De Marneffe et al., 2019; Plank, 2022; Deng et al., 2023).
fnum@section3 A general framework
@@
Domain generalization (DG) can be framed in a similar way. Let $\mathcal{D}$ denote a set of possible domains, where the term domain is a synonym for a joint distribution of $X$ and $Y$. Let $D$ be a random variable on $\mathcal{D}$. Furthermore, let $M$ be a random variable on a space $\mathcal{M}$ that, intuitively, provides partial information about $D$. The idea in DG is that $D$ determines a distribution of $(X,Y)$, but is not observed. $M$ provides partial information (or metadata) about $D$, and is thus useful at test time in adapting the classifier to the test domain.
Remark 1
While the choice of $M$ will depend on the application, one choice that is always viable is to take $M$ to be $P_{X|D}$, the marginal distribution $X$ for the given domain, which is known at test time though the unlabeled test sample.
As we argue below, the observability of $M$ is what makes DG an interesting problem, and distinct from standard classification.
Formally, we assume that $(X,Y,M,D)$ are jointly distributed, with joint distribution denoted $P_{XYMD}$. This distribution induces several other distributions of interest in this paper. We follow convention in denoting marginal distributions by keeping the relevant subscripts. For example, $P_{XYD}$ denotes the joint distribution of $(X,Y,D)$ after $M$ is marginalized out. Similarly, $P_{XY}$ denotes the marginal distribution of $(X,Y)$.
For any fixed $d\in\mathcal{D}$, $P_{XY|D=d}$ is a joint distribution of $(X,Y)$. Note that our notation is somewhat redundant, as both $d$ and $P_{XY|D=d}$ are notations for a domain == a joint distribution of $(X,Y)$, but these two notations both play a role our discussion. [^2] Blanchard et al. (2011, 2021) use $P_{XY}$ to denote a random domain, whereas in our notation, a random domain is either $P_{XY|D}$ or just $D$. Our introduction of $D$ for a random domain allows us to use $P_{XY}$ for the “average” domain, which will be a critical concept in what follows.
@@
This implies that, given $D$, the joint distribution of $X$ and $Y$ does not change with knowledge of $M$. An important special case where this holds is when $M=g(D)$ for some deterministic $g:\mathcal{D}\to\mathcal{M}$. Table 1 lists three applications that illustrate this probabilistic framework, and that will be developed in Section˜6.
@@
Sentiment analysis
(Multiple Annotators)
Sentence
Sentiment label
(e.g., positive, negative)
Annotator identity
(e.g., “Annotator 1”)
Annotator’s demographic profile
(e.g., age)
Review rating prediction
(Multiple Reviewers)
Product review
Numerical rating
(e.g., 1–5 stars)
Reviewer identity
(e.g., “Reviewer 2”)
Unlabeled texts written by the reviewer
$\{X_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{X|D=d}$
Image classification
across styles
Image
Object category label
(e.g., dog, car)
Image style
(e.g., photograph, painting)
Textual description of style
@@
In practice, this risk is estimated by holding out several of the domains, and averaging the test errors on them.
Remark 2
This probabilistic framing of DG generalizes the original formulation of DG by Blanchard et al. (2011). They focus on the special case where $M$ is the marginal distribution of $X$ for the given domain (aka, $M=P_{X|D}$), and focus on the challenges associated to learning from empirical samples of the training and test $X$-marginals.
@@
$\widehat{f}_{\mathrm{DG}}=\underset{f\in\mathcal{F}}{\operatorname{arg\ min\ }}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(y_{ij},f(x_{ij},m_{i})).$ (3)
We are interested in when DI-ERM outperforms pooling ERM. From a theoretical perspective, we compare the optimal risks associated with these two approaches.
fnum@section4 Risk and Bayes risk
To aid in understanding domain generalization, it is helpful to consider DG in relation to two other problem settings. These settings differ only in what information the classifier $f$ has access to. In all cases, the performance measure is the risk
@@
where the argument of $f(\cdot)$ depends on settings. In this section, we compare the risks and Bayes risks associated to these three settings. This sets the state for our theoretical results in the next section.
@@
where, because $f$ does not depend on $D$ or $M$, these variables marginalize out. Therefore, the problem reduces to learning with respect to the marginal distribution of $(X,Y)$, which corresponds to pooling data across domains. The optimal classifier $f^{*}_{\mathrm{pool}}$ is thus the Bayes classifier for the marginal distribution of $(X,Y)$:
@@
$R^{*}_{\mathrm{pool}}$ is the best possible performance of pooling ERM.
Partial Domain Information:   This is the setting of domain generalization. The classifier has access to not only $x$, but also the metadata $m$ that conveys partial information about the true domain $d$. A classifier in this setting is denoted $f(x,m)$. The risk is
@@
The optimal classifier $f^{*}_{\mathrm{DG}}$ is now the Bayes classifier for the distribution of $X,Y|M$,
$f^{*}_{\mathrm{DG}}(x,m)=\underset{k}{\operatorname{arg\ max\ }}\mathbb{P}\left(Y=k|X=x,M=m\right),$
@@
$R^{*}_{\mathrm{DG}}:=$ $\ \mathbb{E}_{X,Y,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{DG}}(X,M)\neq Y}\right]$
@@
$R^{*}_{\mathrm{DG}}$ is the best possible performance DI-ERM. Clearly, $R^{*}_{\mathrm{pool}}\leq R^{*}_{\mathrm{DG}}$, since adding features (here, the metadata) can only decrease the Bayes risk.
@@
Remark 3
DG is very similar to multi-source domain adaptation (DA). The training data available to the learner is the same in both problems. In DA, however, the goal is to achieve $R^{*}_{\mathrm{full}}$ . Essentially, the difference between DG and DA is that DG seeks to optimize the expected risk on a new, random test domain, whereas DA seeks to optimize risk on a specific, fixed test domain. To achieve this goal, it is common in DA to make additional assumptions about how the test domain relates to the training/source domains. In DG, no such assumptions are necessary since the Bayes classifier only uses observable information.
fnum@section5 Risk Comparison
This section develops bounds that relate the three Bayes risks defined in the previous section, revealing settings where domain information is and is not beneficial. The following basic result provides a starting point.
Proposition 1 (Risk Hierarchy)
$R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DG}}\geq R^{*}_{\mathrm{full}}.$
The proof is straightforward (see Section˜C.1). The first inequality is trivial, as extending a feature vector can never decrease the Bayes risk. The second inequality follows from (1).
We next describe distributional assumptions under which these inequalities become strict, and with a quantifiable gap. Under such assumptions, domain information yields improved predictive capability compared to prediction with feature vectors only. Toward that end, consider the following definition.
Definition 1 (Point-wise Margin)
Consider any random triple $(X,Y,M)$, where $Y$ is discrete. Define the point-wise margin of $Y|M=m,X=x$ as,
@@
The operator $\operatorname{2nd}\max_{k}$ returns the second largest value of its argument. Thus, if the two largest values of $\mathbb{P}\left(Y=k|X=x,M=m\right)$ are the same, $\gamma(x,m)=0$. Intuitively, $\gamma(x,m)$ reflects the degree of certainty that the Bayes classifier $f^{*}_{\mathrm{DG}}(x,m)$ has about its prediction. The larger $\gamma(x,m)$, the more confident the prediction.
The next result gives upper and lower bounds on the gap between $R^{*}_{\mathrm{DG}}$ and $R^{*}_{\mathrm{pool}}$. This gap is the additional reduction in risk that results from leveraging the partial domain information $M$.
Theorem 1 (Risk Reduction w/ Domain Info)
@@
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}$
$\quad\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}\right].$
The proof of Theorem 1 is in Section˜C.2. The upper bound represents the probability of disagreement between the domain-informed classifier $f^{*}_{\mathrm{DG}}$ and the pooled classifier $f^{*}_{\mathrm{pool}}$. The lower bound can be interpreted as the expected cost of disagreement, where the cost is zero when the predictions agree, and equals the margin $\gamma(X,M)$ when they differ. Hence, domain information is particularly beneficial when $f^{*}_{\mathrm{DG}}$ frequently disagrees with $f^{*}_{\mathrm{pool}}$ in regions of high confidence. Theorem˜1 also implies necessary and sufficient conditions for the risk gap being zero. If $f^{*}_{\mathrm{DG}}=f^{*}_{\mathrm{pool}}$ almost surely, then $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}=0$. If $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DG}}=0$, then $\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DG}}(X,M)}$ must be zero almost surely, meaning $f^{*}_{\mathrm{DG}}$ and $f^{*}_{\mathrm{pool}}$ can only disagree whenever two or more classes achieve the maximum posterior probability. Figure˜1 gives more intuition.
Figure 1: Illustration of Theorem˜1. Consider binary classification with $X\in\mathbb{R}$, $Y\in\{1,2\}$, and $M\in\{1,2\}$. Then the Bayes classifiers $f^{*}_{\mathrm{pool}}(x)$, $f^{*}_{\mathrm{DG}}(x,m=1)$ and $f^{*}_{\mathrm{DG}}(x,m=2)$ can be obtained by thresholding the corresponding posteriors at $1/2$. The left figure shows a scenario where the domain-informed classifier $f^{*}_{\mathrm{DG}}$ and the pooled classifier $f^{*}_{\mathrm{pool}}$ agree everywhere, and therefore both upper and lower bound are 0. In this case, domain information $M$ is not beneficial. The right figure shows a scenario where $f^{*}_{\mathrm{DG}}$ disagrees with $f^{*}_{\mathrm{pool}}$ in certain regions, and domain information does lead to lower Bayes risk.

### v2

# Domain Generalization Under Posterior Drift

@@ Abstract
Domain generalization (DG) is the problem of generalizing from several distributions (or domains), for which labeled training data are available, to a new test domain for which no labeled data is available. For the prevailing benchmark datasets in DG, there exists a single classifier that performs well across all domains.
In this work, we study a fundamentally different regime where the domains satisfy a posterior drift assumption, in which the optimal classifier might vary substantially with domain. We establish a decision-theoretic framework for DG under posterior drift, and investigate the practical implications of this framework through experiments on language and vision tasks.
1 Introduction
@@
Domain generalization (DG) is the learning problem where the learner has access to labeled data from several source domains, and the goal is to generalize to a new target domain for which no labeled data is available. More precisely, DG assumes labeled training data from each of several training domains $D_{1},\ldots,D_{N}$, and the goal is to produce a classifier that works on any test domain $D$ from which no labeled data are available.
DG was first formulated by Blanchard et al. (2011). In this work and the follow-up work of Muandet et al. (2013), prediction of a label $Y$ is based not just on a feature vector $X$ from a test domain, but also on the marginal distribution $P_{X|D}$ of feature vectors from the test domain . Thus, $P_{X|D}$ allows the predictor to adapt to the test domain.
Most subsequent work on DG, however, has sought to learn a classifier that predicts $Y$ solely from $X$, thus ignoring domain information during inference. This is reflected in recent surveys:
  • Wang et al. (2022b) (survey): “The goal of domain generalization is to learn a robust and generalizable predictive function $h:\mathcal{X}\to\mathcal{Y}$ from the $M$ training domains to achieve a minimum prediction error on an unseen test domain $S_{\text{test}}$.”
This focus on domain agnostic DG, in which domain information is not used at inference, can be traced largely to the datasets employed in DG research. While Blanchard et al. (2011) studied an application in flow cytometry, most DG research since then has focused on applications in computer vision. For many vision tasks in the DG literature, there exists a single classifier that performs well on all domains. In such settings, strong performance is indeed possible without leveraging domain-specific information at test time.
As a consequence, both the theory and methodology of DG remain relatively underdeveloped for settings in which the Bayes-optimal classifier varies substantially across domains. While some recent methods incorporate domain information into training and inference (Dubey et al., 2021; Zhang et al., 2021; Yao et al., 2024), the fundamental role of domain information, namely, its precise decision-theoretic value and the conditions under which it strictly improves performance, remains poorly understood.
To address this limitation, we introduce a theoretical framework for domain-informed DG. This framework is grounded in a novel statistical formulation of DG that extends both the original formulation introduced by Blanchard et al. (2011) and the domain-agnostic setting. It allows that domain metadata — features that describe the domain — is available during training and inference.
Leveraging this framework, we establish risk bounds that demonstrate the benefit of domain information under posterior drift, a new class of DG problems where the optimal classifier changes substantially with the domain (Theorems 1 and 2). In particular, we show that under posterior drift, domain-agnostic methods are provably suboptimal. We further provide a condition, reflecting vision tasks as described above, under which domain-agnostic methods are sufficient.
To demonstrate the practical implications of our theory, we experimentally investigate the performance of domain-informed empirical risk minimization (DI-ERM), the natural ERM procedure induced by our formulation, and compare it to leading DG methods on language and vision tasks.
Our contributions are summarized as follows:
  • A statistical formulation of DG that incorporates domain metadata and generalizes the standard DG setting.
  • Risk bounds characterizing the decision-theoretic value of domain metadata, particularly under a newly identified posterior drift class of DG problems.
  • A formal condition under which domain-agnostic DG is sufficient, providing a theoretical explanation for the widely observed phenomenon that domain-agnostic ERM is “hard to beat” on standard benchmarks.
  • A quantification of the decision-theoretic gap between DG and domain adaptation, resolving an open question posed in Blanchard et al. (2021, Lemma 9).
  • Empirical validation of our theory on both language and vision benchmarks.
2 Literature Review
In these vision-based setups, while the $X$-marginal changes across the domains, a single function of the input $X$ realizes the Bayes classifier for all domains. Under this regime, domain-specific features are often regarded as irrelevant or even spurious (Sagawa et al., 2020; Bai et al., 2025) for predicting labels. Consequently, much of the literature has focused on learning domain-invariant representations. (Sun and Saenko, 2016; Ganin et al., 2016; Arjovsky et al., 2019). Additional references are in Appendix A.
@@ 2 Literature Review
Despite extensive efforts to design advanced DG algorithms, a persistent puzzle is the surprising effectiveness of what we will call pooling empirical risk minimization (pooling-ERM), a baseline that simply pools labeled data from all source domains together and trains a domain-agnostic classifier. Multiple studies have consistently found that pooling-ERM is hard to beat:
@@
Similar observations have been made on other DG benchmarks (Koh et al., 2021; Sagawa et al., 2022), as well as in related settings such as federated domain generalization (Bai et al., 2024).
We argue that the key to understanding these observations lies in the probabilistic assumptions governing distributional shift across domains. Different from prior works, our work examines an assumption called posterior drift, previously studied in domain adaptation (Scott, 2019; Cai and Wei, 2021; Maity et al., 2024; Zhu et al., 2024; Wang and Ning, 2025), where the conditional distribution of $Y|X$ varies across domains. This type of shift commonly arises in natural language processing (NLP). For a given sentence $X$, different annotators (or populations) may interpret its semantic content differently, leading to divergent labels $Y$ (e.g., offensive vs. non-offensive, positive vs. negative). This form of ambiguity and annotator disagreement has been documented across a wide range of NLP applications (De Marneffe et al., 2019; Plank, 2022; Deng et al., 2023).
In this work, we formulate posterior drift in a DG context, and develop statistical theory for DG problems characterized by posterior drift. Under this regime, pooling-ERM is provably suboptimal, and strictly stronger performance is achievable by domain-informed ERM (defined in the next section). Our theory highlights the value of using domain metadata to perform domain adaptive prediction. Through this lens, Yao et al. (2024) is closest to our work. They leverage domain metadata to re-weight prediction during inference time.
3 A General Statistical Framework
@@
Domain generalization (DG) can be framed in a similar way. Let $\mathcal{D}$ denote a set of possible domains, where the term domain is a synonym for a joint distribution of $X$ and $Y$. Let $D$ be a random variable on $\mathcal{D}$. Furthermore, let $M$ be a random variable on a space $\mathcal{M}$ that, intuitively, provides partial information about $D$. In DG, $D$ determines a distribution of $(X,Y)$, but is not observed. $M$ provides partial information (or metadata) about $D$, and is thus useful at test time in adapting the classifier to the test domain.
Remark 1.
While the availability of metadata $M$ will depend on the application, one choice that is frequently viable is to take $M=P_{X|D}$, the marginal distribution $X$ for the given domain, which is known at test time through the unlabeled test sample.
Below we argue that the observability of $M$ is what makes DG distinct from standard classification in a certain decision-theoretic sense. Formally, we assume that $(X,Y,M,D)$ are jointly distributed, with joint distribution denoted $P_{XYMD}$. This distribution induces several other distributions of interest in this paper. We follow convention in denoting marginal distributions by keeping the relevant subscripts. For example, $P_{XYD}$ denotes the joint distribution of $(X,Y,D)$ after $M$ is marginalized out. Similarly, $P_{XY}$ denotes the marginal distribution of $(X,Y)$.
For any fixed $d\in\mathcal{D}$, $P_{XY|D=d}$ is a joint distribution of $(X,Y)$. Note that our notation is somewhat redundant, as both $d$ and $P_{XY|D=d}$ refer to a domain, namely, a joint distribution of $(X,Y)$, but these two notations both play a role in our discussion. [^2] Blanchard et al. (2011, 2021) use $P_{XY}$ to denote a random domain, whereas in our notation, a random domain is either $P_{XY|D}$ or just $D$. Our introduction of $D$ for a random domain allows us to use $P_{XY}$ for the “average” domain, which will be a critical concept in what follows.
@@
This implies that, given $D$, the joint distribution of $X$ and $Y$ does not change with knowledge of $M$. An important special case where this holds is when $M=g(D)$ for some deterministic $g:\mathcal{D}\to\mathcal{M}$. Table 1 lists several applications that illustrate this probabilistic framework, and will be empirically evaluated in Section 6.
@@
Sentiment analysis (Multiple Annotators) Sentence Sentiment label (e.g., positive, negative) Annotator identity (e.g., “Annotator 1”) Annotator’s demographic profile (e.g., age)
Review rating prediction (Multiple Reviewers) Product review Numerical rating (e.g., 1–5 stars) Reviewer identity (e.g., “Reviewer 2”) Unlabeled texts written by the reviewer $\{X_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}P_{X|D=d}$
Image classification across styles Image Object category label (e.g., dog, car) Image style (e.g., photograph, painting) Textual description of style
@@
where the argument of $f(\cdot)$ depends on settings. In practice, this risk is estimated by holding out several of the domains, and averaging the test errors on them.
Remark 2.
This probabilistic framing of DG generalizes the original formulation of DG by Blanchard et al. (2011). They focus on the special case where $M$ is the marginal distribution of $X$ for the given domain ($M=P_{X|D}$), and focus on the challenges associated with learning from empirical samples of the training and testing $X$-marginals.
@@
$\widehat{f}_{\mathrm{DI}}=\underset{f\in\mathcal{F}}{\operatorname{arg\ min\ }}\frac{1}{N}\sum_{i=1}^{N}\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\ell(y_{ij},f(x_{ij},m_{i})).$ (3)
To highlight the importance of using domain metadata, we are interested in when DI-ERM outperforms pooling-ERM. In the next section, we compare the optimal risks associated with these two approaches.
4 Risk and Bayes Risk
To aid in understanding domain generalization, it is helpful to understand the risk and Bayes risk depending on the information available to the classifier. In this section, we examine the risk
@@ 4 Risk and Bayes Risk
where the argument of $f(\cdot)$ is determined by one of three possible settings. This sets the stage for our main results in the next section.
@@
where, because $f$ does not depend on $D$ or $M$, these variables marginalize out. Therefore, the problem reduces to learning with respect to the marginal distribution of $(X,Y)$, which can be viewed as the “average” domain. The optimal classifier $f^{*}_{\mathrm{pool}}$ is thus the Bayes classifier for the marginal distribution of $(X,Y)$:
@@
$R^{*}_{\mathrm{pool}}$ is the best possible performance of pooling-ERM.
Partial Domain Information:   This is our setting of domain generalization. The classifier has access to not only $x$, but also the metadata $m$ that conveys partial information about the true domain $d$. A classifier in this setting is denoted $f(x,m)$. The risk is
@@
The optimal “domain-informed” classifier $f^{*}_{\mathrm{DI}}$ is now the Bayes classifier for the distribution of $X,Y|M$,
$f^{*}_{\mathrm{DI}}(x,m)=\underset{k}{\operatorname{arg\ max\ }}\mathbb{P}\left(Y=k|X=x,M=m\right),$
@@
$R^{*}_{\mathrm{DI}}:=$ $\ \mathbb{E}_{X,Y,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{DI}}(X,M)\neq Y}\right]$
@@
$R^{*}_{\mathrm{DI}}$ is the best possible performance of DI-ERM. Clearly, $R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DI}}$, since adding features (here, the metadata) can only decrease the Bayes risk.
@@
Remark 3.
DG is very similar to multi-source domain adaptation (DA). The training data available to the learner is the same in both problems. In DA, however, the performance metric of interest is $R^{*}_{\mathrm{full}}$. Essentially, the difference between DG and DA is that DG seeks to optimize the expected risk on a new, random test domain, whereas DA seeks to optimize risk on a specific, fixed test domain.
5 Risk Comparison
This section develops bounds that relate the three Bayes risks defined in the previous section, identifying when domain information is beneficial and when it is not. The following basic result provides a starting point.
Proposition 1 (Risk Hierarchy).
$R^{*}_{\mathrm{pool}}\geq R^{*}_{\mathrm{DI}}\geq R^{*}_{\mathrm{full}}.$
The proof is straightforward (see Section C.1). The first inequality holds as adding more feature can never decrease the Bayes risk. The second inequality follows from Assumption (1).
Our goal in this section is to describe distributional assumptions under which these inequalities become strict, and with a quantifiable gap. Toward that end, consider the following.
Definition 1 (Point-wise Margin).
Consider any random triple $(X,Y,M)$, where $Y$ is discrete. Define the point-wise margin of $Y|X=x,M=m$ as,
@@ 5 Risk Comparison
The operator $\operatorname{2nd}\max_{k}$ returns the second largest value of its argument. Thus, if the two largest values of $\mathbb{P}\left(Y=k|X=x,M=m\right)$ are the same, $\gamma(x,m)=0$. Intuitively, $\gamma(x,m)$ reflects the degree of certainty that the Bayes classifier $f^{*}_{\mathrm{DI}}(x,m)$ has about its prediction. The larger $\gamma(x,m)$, the more confident the prediction.
The next result gives upper and lower bounds on the gap between $R^{*}_{\mathrm{DI}}$ and $R^{*}_{\mathrm{pool}}$. This gap is the additional reduction in risk that results from leveraging the partial domain information $M$.
Theorem 1 (Risk Reduction w/ Domain Info).
@@
$\mathbb{E}_{X,M}\left[\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right]$
$\quad\leq R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}$
$\quad\leq\mathbb{E}_{X,M}\left[\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}\right].$
The proof of Theorem 1 is in Section C.2. The upper bound represents the probability of disagreement between the domain-informed classifier $f^{*}_{\mathrm{DI}}$ and the pooling classifier $f^{*}_{\mathrm{pool}}$. The lower bound can be interpreted as the expected cost of disagreement, where the cost is zero when the predictions agree, and equals to the margin $\gamma(X,M)$ when they differ. Hence, domain information is particularly beneficial when $f^{*}_{\mathrm{DI}}$ frequently disagrees with $f^{*}_{\mathrm{pool}}$ in regions of high confidence. Theorem 1 also implies necessary and sufficient conditions for the risk gap being zero. If $f^{*}_{\mathrm{DI}}=f^{*}_{\mathrm{pool}}$ almost surely, then $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}=0$. If $R^{*}_{\mathrm{pool}}-R^{*}_{\mathrm{DI}}=0$, then $\gamma(X,M)\mathbbm{1}_{f^{*}_{\mathrm{pool}}(X)\neq f^{*}_{\mathrm{DI}}(X,M)}$ must be zero almost surely, meaning $f^{*}_{\mathrm{DI}}$ and $f^{*}_{\mathrm{pool}}$ can only disagree whenever two or more classes achieve the maximum posterior probability. Figure 1 gives more intuition.
Figure 1: Illustration of Theorem 1. Consider binary classification with $X\in\mathbb{R}$, $Y\in\{1,2\}$, and $M\in\{1,2\}$. Then the Bayes classifiers $f^{*}_{\mathrm{pool}}(x)$, $f^{*}_{\mathrm{DI}}(x,m=1)$ and $f^{*}_{\mathrm{DI}}(x,m=2)$ can be obtained by thresholding the corresponding posteriors at $1/2$. The left figure shows a scenario where the domain-informed classifier $f^{*}_{\mathrm{DI}}$ and the pooled classifier $f^{*}_{\mathrm{pool}}$ agree everywhere, and therefore both upper and lower bound are 0. In this case, domain information $M$ is not beneficial. The right figure shows a scenario where $f^{*}_{\mathrm{DI}}$ disagrees with $f^{*}_{\mathrm{pool}}$ in certain regions, and domain information does lead to lower Bayes risk.